Need help ASAP
Price-sales raio MacBook Air
Answer #1

a)

Net income = retained earnings + dividends

Net income = 643,000 + 40,000 = $683,000

Earnings per share = Net income / shares

earnings per share = 683,000 / 750,000

earnings per share = $0.91

b)

Dividends per share = dividends / shares

Dividends per share = 40,000 / 750,000

Dividends per share = $0.05

c)

Please not: The figure of total equity is not very clear. i am assuming it to be $7.38 million. If it has a different figure, please put it in comments and i will re do the question. This will affect market-to-book ratio as well

Book value per share = Total equity / shares outstanding

Book value per share = 7,380,000 / 750,000

Book value per share = $9.84

d)

Market-to-book ratio = Market price / book price

Market-to-book ratio = 30.80 / 9.84

Market-to-book ratio = 3.13

e)

sales per share = sales / shares

sales per share = 10,680,000 / 750,000

sales per share = $14.24

Price-earnings ratio = Price / sales per share

Price-earnings ratio = 30.80 / 14.24

Price-earnings ratio = 2.16